# Deep Learning-based Intrusion Detection Systems: A Survey

### v1

@@ 7.1.3. Potential Heavy Computation Requirements
Similar to DL techniques, DL-IDS also requires a large amount of computing resources to improve their performance. According to (Roberts et al., 2024), the generalizability of neural models is proportional to the investment of computing resources. To illustrate, we report the computing resources of some well-known DL-IDS in Table 7. It shows that existing DL-IDS are all relatively lightweight, as manifested by only one NVIDIA Tesla V100 GPU or only one NVIDIA GeForce RTX 6000 GPU. However, the requirements for computing resources are related to the number of log data. When the challenge of insufficient data is mitigated and more data is available, more computing resources are inevitably required. Meanwhile, we will illustrate in Section 7.2 that there are plenty of powerful techniques that have not been introduced in DL-IDS, which similarly will bring in computation requirements.
Additionally, several DL-IDS (Han et al., 2020; Cheng et al., 2024; Wu et al., 2022; Trivedi et al., 2023; Duan et al., 2021; Meng et al., 2019) have been specifically engineered for real-time intrusion detection. For example, the experiments in Unicorn (Han et al., 2020) demonstrate that the processing time of Unicorn is linear to its workloads with only one core. Yet the computational units, either CPU or GPU, all have their performance upper bound. Unicorn should be implemented in parallel if its workload is too high.
@@ 7.1.4. Future Directions
To conclude, the challenges for DL-IDS in fundamental resources consist of data quality, data volume, and computational overhead. Apart from unintentional errors and non-technical issues in fundamental resources, the research questions that urgently need to be addressed include the contradiction between unaffordable manual labeling and non-generalizable auto-labeling techniques, non-unified benchmark datasets and performance metrics, as well as potential heavy computational overheads. Therefore, we summarize the future directions as follows:
@@ 7.2. Pre-training Theories and Techniques
With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023). Nevertheless, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 8. We can observe that LLMs developed at full speed beginning in 2019. Their prosperity, however, has not extended to DL-IDS. Until now, the only two DL-IDS that incorporate pre-training techniques, AirTag (Ding et al., 2023b) and MAGIC (Jia et al., 2024), still do not make full use of the potential of LLMs. AirTag pre-trains a BERT model on application logs and detects intrusions in terms of embeddings generated by BERT. MAGIC introduces GraphMAE (Hou et al., 2022), a model architecture derived from Graph Autoencoder (Kipf and Welling, 2016) in 2016 but integrated with the famous masked self-supervised learning method (He et al., 2022) in 2022, to conduct self-supervised learning on provenance graphs. MAGIC further designs an adapter to apply the pre-trained model in different detection scenarios. Nevertheless, both AirTag and MAGIC can be regarded as preliminary explorations of pre-training techniques. According to the scaling law (Kaplan et al., 2020), the performance of LLMs will steadily improve, as the parameters, data, and computation increase. And the reasoning ability of LLMs will suddenly emerge (Wei et al., 2022), allowing them to chat with humans smoothly. Such advantageous abilities obviously have not been incorporated into DL-IDS.
@@
Figure 8. Interactions between DL models and DL-IDS. While DL models proposed before 2019 have already leveraged in DL-IDS, the emerging LLMs since 2020 remains underdeveloped in this domain.
@@ 7.2.2. Short of Statistical Log Modeling
LLMs are developed on the basis of statistical language modeling (Rosenfeld, 2000; Jozefowicz et al., 2016), which is not insufficiently studied for log data. The statistical modeling of natural language can be traced back to the early 1950s, when Shannon pioneered the technique of predicting the next element of natural language text (Shannon, 1948) and discussed the n-gram model for English (Shannon, 1951). After that, as machine learning came into view of the NLP research communities, language modeling flourished, and many models such as TreeBank (Marcus et al., 1993), word2vec (Mikolov et al., 2013a, b) and LSTM (Hochreiter and Schmidhuber, 1997) were proposed. Over decades, researchers in NLP have gained solid knowledge of language modeling, whose interests gradually shifted to efficiency. An epoch-making model, Transformer (Vaswani et al., 2017), was presented using the multi-head self-attention mechanism to fulfill parallel computing, which was widely exploited in popular pre-trained models such as BERT (Devlin et al., 2019) and GPT (Achiam et al., 2023) afterward. It is evident that the success of LLMs comes from the prolonged studies on statistical language modeling.
@@
Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped. By contrast, as illustrated in Table 8, other data types whose statistical modeling studies started earlier have already done well in pre-training techniques. For example, Hindle et al. (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language. Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023c, 2024; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection make great progress, many of which are proven to be greatly strengthened by pre-training techniques. Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
Log data is also created by humans, similar to text, speech, and source code. It is generated according to developer-defined log templates, with a form of natural language (e.g., application logs) and formal language (e.g., data provenance in CDM format). Given the fact that natural language (e.g., text and speech) and formal language (e.g., source code) both exhibit positive features in statistical modeling, log data urgently demands similar achievements to facilitate its pre-training research. Although several works (Michael et al., 2020; Inam et al., 2023) have discussed the features of log data, they are orthogonal to the explainable combination of DL and IDS. Compared with the other data types, challenges in statistical log modeling, for instance, may lie in that logs are extremely long and detailed for reliable purposes. It is very common that the length of one single log entry is the same as that of one paragraph in natural language texts. This happens to be the shortcomings of LLMs - not being trustworthy in generated contents and not being able to handle long text.
@@ 7.2.3. Future Directions
Future Directions Investigating how and where to introduce LLMs into DL-IDS, with the objective of balancing the generalizability provided by LLMs and the reliability required by DL-IDS. Exploring fundamental statistical modeling theories for log data. On this basis, designing pre-training frameworks for log data and its downstream tasks such as steps within the workflow of DL-IDS (see Section 3.2).
@@ 7.3. Comprehensive Applications and Scenarios
DL-IDS possess abilities that the traditional IDS lack, or are difficult to realize, such as generalizability for zero-day attacks and modeling ability for complicated downstream tasks. We will elaborate the possible new-style applications and discuss the challenges in and introduced by them.
@@ 7.3.1. Limited Forward and Backward Tracing Scope
In this way, the detection of DL-IDS has been extended to the broadly defined intrusions, including their attributes such as root causes and control measures. When executing backward tracing analysis, DL-IDS are not only required to detect the starting symptom nodes, but also required to find the root causes of these nodes (or vulnerabilities in source codes). In the forward tracing analysis, except for detecting the symptom nodes affected by intrusions, DL-IDS should perform an in-depth analysis to discover the potential symptom nodes and provide control measures for the intrusions.
Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019). In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis. Root causes of code vulnerabilities are thus allowed to derive from intrusions. Pedro et al. (Manso et al., 2019) investigate detection and mitigation for DDoS attacks, which can be regarded as control measures against specific intrusions. We note that these research attempts are all based on heuristics and can only achieve limited efficacy, either in generating root causes using pre-defined rules, or in developing control measures for specific intrusions. The main obstacle is the lack of robust modeling and reasoning capabilities, which gives rise to a substantial need for the introduction of advanced techniques like DL.
@@ 7.3.2. Concerns about Data-driven Adversarial Attacks
To clarify the detection performance, DL-IDS commonly idealize the parts other than detection models in their threat model, of which the most commonly idealized part is data. Such idealization, however, leaves DL-IDS with weaknesses that could be exploited by invaders.
One usual assumption for data is related to data poisoning attacks, as (Rehman et al., 2024; Hassan et al., 2019; Jia et al., 2024; Han et al., 2020) did. In this assumption, no attacks are considered to compromise the security of the log collection systems, that is, log data utilized in DL-IDS is absolutely harmless. But apparently, as attacks become more stealthy and complicated, it is impossible to satisfy such an assumption to some extent. Besides, the harmful data could emerge through non-attacks. Security analysts may fail to label log entries correctly, resulting in an effect similar to the data poison attacks. Our manual checking empirically identifies that the labeling in threaTrace (Wang et al., 2022a) is not fully consistent with the official ground truth documentation. In many cases, by emulating benign system behaviors, the attackers are able to plant backdoors in the training data for DL-IDS, putting the computer systems at invading risk.
The robustness of DL-IDS is also challenged by evasion attacks. To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attack), making them hard to be detected. By early 2002, David et al. (Wagner and Soto, 2002) had indicated the danger of mimicry attacks on HIDS. Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection. While recognized to be powerful, DL-IDS are plagued by the issue that even a trivial perturbation in the data can result in a successful evasion attack (Chakraborty et al., 2018). Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks. However, as noted in recent work (Mukherjee et al., 2023; Goyal et al., 2023, 2024), data-driven attacks still remain a major challenge for DL-IDS.
@@ 7.3.3. Underexplored Promising Scenarios
DL-IDS show excellent performance in the protection of computer and network systems recently. Unfortunately, there are still many promising scenarios for DL-IDS that have not been explored sufficiently yet.

### v2

@@ 7.1.3. Potential Heavy Computation Requirements
Similar to other DL techniques, DL-IDS also requires a potentially large amount of computing resources to improve their performance. According to (Roberts et al., 2024), the generalizability of neural models is proportional to the investment of computing resources. Supposing that the challenge of insufficient data is mitigated and a large volume of log data is available, more computing resources are inevitably required. Besides, we will illustrate in Section 7.2 that there are plenty of powerful techniques that have not been introduced in DL-IDS, which will also bring in computation requirements. Unfortunately, acceleration methods like parallel computation and efficient retrieval have not been fully scheduled by the cybersecurity community. An example is that the computation time of Unicorn equipped with one core is proven linear to its workloads (Han et al., 2020). It is clear that the efficiency of Unicorn, which is not implemented in parallel, will reach the bottleneck as this core does.
@@ 7.1.4. Future Directions
To conclude, the challenges for DL-IDS in fundamental resources consist of data quality, data volume, and computational overhead. Apart from unintentional errors and non-technical issues in fundamental resources, the research questions that urgently need to be addressed include the contradiction between unaffordable manual labeling and non-generalizable auto-labeling techniques, non-unified benchmark datasets and evaluation metrics, as well as potential heavy computational overheads. Therefore, we summarize the future directions as follows:
@@ 7.2. Pre-training Theories and Techniques
With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024b), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023). In contrast, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 6. We can observe that LLMs developed at full speed beginning in 2019. Their prosperity, however, has not extended to DL-IDS. Until now, the only two DL-IDS that incorporate pre-training techniques, AirTag (Ding et al., 2023b) and MAGIC (Jia et al., 2024), still do not make full use of the potential of LLMs. AirTag pre-trains a BERT model on application logs and detects intrusions in terms of embeddings generated by BERT. MAGIC introduces GraphMAE (Hou et al., 2022), a model architecture derived from Graph Autoencoder (Kipf and Welling, 2016) in 2016 but integrated with the famous masked self-supervised learning method (He et al., 2022) in 2022, to conduct self-supervised learning on provenance graphs. MAGIC further designs an adapter to apply the pre-trained model in different detection scenarios. Nevertheless, both AirTag and MAGIC can be regarded as preliminary explorations of pre-training techniques. According to the scaling law (Kaplan et al., 2020), the performance of LLMs will steadily improve, as the parameters, data, and computation increase. And the reasoning ability of LLMs will suddenly emerge (Wei et al., 2022), allowing them to chat with humans smoothly. Such advantageous abilities obviously have not been incorporated into DL-IDS.
@@
Figure 6. Interactions between DL models and DL-IDS. While DL models proposed before 2019 have already leveraged in DL-IDS, the emerging LLMs (or pre-training theories and thechniques) since 2020 remains underdeveloped in this domain.
@@ 7.2.2. Short of Statistical Log Modeling
LLMs are developed on the basis of statistical language modeling (Rosenfeld, 2000; Jozefowicz et al., 2016), which is not insufficiently studied for log data. The statistical modeling of natural language can be traced back to the early 1950s when Shannon pioneered the technique of predicting the next element of natural language text (Shannon, 1948) and discussed the n-gram model for English (Shannon, 1951). After that, as machine learning came into view of the NLP research communities, language modeling flourished, and many models such as TreeBank (Marcus et al., 1993), word2vec (Mikolov et al., 2013a, b) and LSTM (Hochreiter and Schmidhuber, 1997) were proposed. Over decades, researchers in NLP have gained solid knowledge of language modeling, whose interests gradually shifted to efficiency. An epoch-making model, Transformer (Vaswani et al., 2017), was presented using the multi-head self-attention mechanism to fulfill parallel computing, which was widely exploited in popular pre-trained models such as BERT (Devlin et al., 2019) and GPT (Achiam et al., 2023) afterward. It is evident that the success of LLMs comes from the prolonged studies on statistical language modeling.
@@
Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped. By contrast, as illustrated in Table 7, the statistical modeling studies of other types of data have already started. Hindle et al. (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language. Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs. Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
We argue that log data is also created by humans, similar to text, speech, and source code. It is generated according to developer-defined log templates, with a form of both natural language (e.g., application logs) and formal language (e.g., data provenance in CDM format). Given the fact that natural language (e.g., text and speech) and formal language (e.g., source code) both exhibit positive performance in pre-training, log data urgently demands statistical modeling achievements to facilitate its pre-training research. Although several works (Michael et al., 2020; Inam et al., 2023) have discussed the features of log data, they are orthogonal to the explainable combination of DL and IDS. Compared with the other data types, challenges in statistical log modeling, for instance, may lie in that logs are extremely long and detailed for reliable purposes. It is very common that the length of one single log entry is the same as that of one paragraph in natural language texts. These challenges happen to be the shortcomings of LLMs - not being able to handle long text and not being trustworthy in generated contents.
@@ 7.2.3. Future Directions
Future Directions Investigating how and where to introduce LLMs into DL-IDS like (Ni et al., 2025), with the objective of balancing the generalizability provided by LLMs and the reliability required by DL-IDS. Exploring fundamental statistical modeling theories for log data. On this basis, designing pre-training frameworks for log data and its downstream tasks such as steps within the workflow of DL-IDS (see Section 3.2).
@@ 7.3. Comprehensive Applications and Scenarios
DL-IDS possess abilities that the traditional IDS lack, or are difficult to realize, such as generalizability for zero-day attacks and modeling ability for complicated downstream tasks. We will elaborate on the possible new-style applications and discuss the challenges in and introduced by them.
@@ 7.3.1. Limited Forward and Backward Tracing Scope
In this way, the detection of DL-IDS has been extended to the broadly defined intrusions, including their attributes of both root causes and control measures. When executing backward tracing analysis, DL-IDS are not only required to detect the starting symptom nodes of intrusions, but also required to find the root causes of these symptom nodes (i.e., vulnerabilities in source codes). In the forward tracing analysis, except for detecting the symptom nodes affected by intrusions, DL-IDS should perform an in-depth analysis to discover the potentially compromised nodes and provide control measures for handling intrusions.
Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019). In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis. Root causes for the exploited vulnerabilities are capable of directly deriving from intrusion detection. Pedro et al. (Manso et al., 2019) investigate detection and mitigation methods for DDoS attacks, aiming to control them immediately. We note that these research attempts are all based on heuristics, either using pre-defined rules to generate root causes, or developing control measures for specific intrusions. Thus, there is a substantial need to introduce advanced DL techniques to this problem.
@@ 7.3.2. Concerns about Data-driven Adversarial Attacks
To validate the detection performance, DL-IDS commonly idealize the experimental data in their threat model. Such idealization, however, leaves DL-IDS with weaknesses that could be exploited by invaders. For example, a common assumption is that no attacks are considered to compromise the security of the log collection systems (Rehman et al., 2024; Hassan et al., 2019; Jia et al., 2024; Han et al., 2020), namely log data utilized in DL-IDS is absolutely harmless. But as attacks become more stealthy and complicated, it is impossible to satisfy such an assumption apparently. When DL-IDS encounter intentional data poisoning attacks, prediction backdoors could be easily planted as persistent vulnerabilities.
The robustness of DL-IDS is also challenged by data-driven evasion attacks. To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attacks), making them hard to be detected. By early 2002, David et al. (Wagner and Soto, 2002) have indicated the danger of mimicry attacks on HIDS. Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection. From a study (Chakraborty et al., 2018), DL-IDS can be even plagued by a trivial perturbation in log data. Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks. However, as noted in recent work (Mukherjee et al., 2023; Goyal et al., 2023, 2024), data-driven attacks still remain a major challenge for DL-IDS.
@@ 7.3.3. Underexplored Promising Scenarios
While DL-IDS show excellent performance in the protection of computer and network systems recently, there are still many promising scenarios for DL-IDS that have not been explored sufficiently.